Determine whether the flow is laminar or turbulent if glycerine at 25degC flows in a pipe with 150mm diameter. The average velocity of flow is 3.6m/s. (mass density = 1259 kg/m3; dynamic viscosity: 9.60 x 10^-1 Pa.s)


1
Expert's answer
2021-12-13T16:55:32-0500

We should find Reynolds number

"Re = \frac{ρVd}{\mu}"

If Re ≤ 2000 the flow is laminar

If Re ≥ 4000 the flow is turbulent

If 2000< Re < 4000 the flow is in transition

Density ρ=1259 kg/m3

Viscosity "\mu = 9.6 \times 10^{-1} \; Pa \cdot sec"

Diameter d = 150 mm = 0.15 m

Velosity = 3.6 m/sec

"Re = \frac{1259 \times 3.6 \times 0.15}{9.6 \times 10^{-1}} = 708.18"

Re<2000

The flow is laminar.